# Parallel Diffusion Solver via Residual Dirichlet Policy Optimization

### v1

@@ IV-C Text-to-Image Experiment Results
LogSNR [lu2022dpm] 54.07 8.88 7.95 3.97
Polynomial [karras2022elucidating] 11.10 8.89 4.50 3.72
@@
Heun [karras2022elucidating] 15.91 6.65 4.61 3.57
iPNDM [liupseudo, krizhevsky2009learning] 13.69 6.64 4.59 3.59
DPM-Solver-2 [lu2022dpm] 10.40 4.33 2.82 2.49
@@ IV-E Ablation Studies
Effect of scaling factors. [ningelucidating, li2024alleviating] identify exposure bias—i.e., the input mismatch between training and sampling—as a key factor leading to error accumulation and sampling drift. To mitigate the bias, they propose scaling the gradient and shifting the timestep. Building on these insights, our EPD-Solver introduces two learnable parameters: $o_{n}$ and $\delta_{n}^{k}$. We compare FID scores without these scaling factors to assess their impact. As shown in  Table IX, omitting the scaling factors noticeably reduces image quality. For instance, without $o_{n}$, FID rises from 4.33 to 5.84 at Para. NFE = 5.
@@
Effect of Residual Dirichlet Policy Optimization. To validate the effectiveness of our parameter-efficient RL fine-tuning, we compare the performance of EPD-Solver before (Stage 1) and after (Stage 2) the Residual Dirichlet Policy Optimization. As illustrated in Figure 9, Stage 2 yields consistent improvements across multiple human alignment metrics compared to the distilled baseline. For instance, on Stable Diffusion v1.5, the RL fine-tuning significantly boosts the ImageReward score from 3.0405 to 3.11212. Similarly, on SD3-Medium ($512\times 512$) at 20 NFE, our Stage 2 solver achieves an HPSv2.1 score of 0.2742, effectively bridging the gap to high-step baselines3. These results confirm that while Stage 1 provides a robust initialization by capturing the trajectory curvature, Stage 2 is crucial for aligning the sampling behavior with human perceptual preferences without increasing inference cost.
@@ A-B Implementation Details of EPD-Plugin
The EPD-Plugin serves as a module integrated in any existing ODE solver. We illustrate this using the multi-step iPNDM [liupseudo, zhangfast] sampler as a representative implementation. We begin with a brief review of the iPNDM sampler.
@@
TABLE X: FID results on the choices of two intermediate points. Evaluations are conducted on CIFAR-10 [krizhevsky2009learning]. Start point: $t_{n+1}$, end point: $t_{n}$, midpoints: $\sqrt{t_{n}t_{n+1}},\frac{1}{2}(t_{n}+t_{n+1})$, and ‘random’ denotes a midpoint randomly chosen from $t_{n},t_{n+1}]$.
@@ A-C Implementation Details of ParaDiGMS
For direct comparison with EDP-{Solver, Plugin}, we re-implemented the ParaDiGMS sampler [shih2023parallel] in the EDM [karras2022elucidating] framework, as its public implementation [^1] https://github.com/AndyShih12/paradigms is tailored for Stable Diffusion. To ensure a fair latency comparison with our single-GPU EPD-Solver, we run ParaDiGMS on two NVIDIA 4090 GPUs, distributing the workload evenly by matching the Para. NFE/GPU ratio.
@@ A-D Further Details on the Text-to-Image Experimental Setup
Figure 13: Analysis on local sampling trajectory. The figure shows the generation path of two randomly selected pixels in the images. We employ the EPD ($\text{Para. NFE}=5,K=2$) sampler for sampling, using the trajectory of its teacher sampler as the target trajectory. We present the sampling trajectories with $\text{NFE}=5$ of DDIM [songdenoising], DPM-Solver [lu2022dpm], and iPNDM [zhangfast] on CIFAR-10 [krizhevsky2009learning].
Figure 14: Comparison of generated samples among DPM-Solver-2 [lu2022dpm], iPNDM [zhangfast] and EPD-Solver. Compared to other samplers, EPD-Solver achieves high-quality results even at NFE = 3. Additional visualizations are provided in section B-C.
@@ A-E Qualitative Analyses of validation experiments
Qualitative results on trajectory. Since visualizing the trajectories of high-dimensional data is challenging, we adopt the analysis framework in [liupseudo]. Specifically, as shown in  fig. 13, we randomly select two pixels from the images to perform local trajectory visualization, illustrating how their values evolve during the sampling process. Given the sampling ${\mathbf{x}}_{t_{N}},{\mathbf{x}}_{t_{N-1}},\dots,{\mathbf{x}}_{t_{0}}$, we track the corresponding values $v_{t}^{1}$ and $v_{t}^{2}$ at two randomly chosen positions $p_{1}$ and $p_{2}$. We then represent $(v_{t}^{1},v_{t}^{2})$ as data points and visualize them in $\mathbb{R}^{2}$. We can clearly observe that the pixel value trajectories of EPD-Solver ($\text{Para. NFE}=5,K=2$) are closer to the target trajectories compared to other samplers. This shows that our EPD-Solver can generate more accurate trajectory, significantly reducing errors in the sampling process.

### v2

@@ IV-C Text-to-Image Experiment Results
LogSNR [34] 54.07 8.88 7.95 3.97
Polynomial [15] 11.10 8.89 4.50 3.72
@@
Heun [15] 15.91 6.65 4.61 3.57
iPNDM [32, 22] 13.69 6.64 4.59 3.59
DPM-Solver-2 [34] 10.40 4.33 2.82 2.49
@@ IV-E Ablation Studies
Effect of scaling factors. [42, 25] identify exposure bias—i.e., the input mismatch between training and sampling—as a key factor leading to error accumulation and sampling drift. To mitigate the bias, they propose scaling the gradient and shifting the timestep. Building on these insights, our EPD-Solver introduces two learnable parameters: $o_{n}$ and $\delta_{n}^{k}$. We compare FID scores without these scaling factors to assess their impact. As shown in  Table IX, omitting the scaling factors noticeably reduces image quality. For instance, without $o_{n}$, FID rises from 4.33 to 5.84 at Para. NFE = 5.
@@
Effect of Residual Dirichlet Policy Optimization. To validate the effectiveness of our parameter-efficient RL fine-tuning, we compare the performance of EPD-Solver before (Stage 1) and after (Stage 2) the Residual Dirichlet Policy Optimization. As illustrated in Figure 9, Stage 2 yields consistent improvements across multiple human alignment metrics compared to the distilled baseline. For instance, on Stable Diffusion v1.5, the RL fine-tuning significantly boosts the ImageReward score from -0.002 to 0.012. Similarly, on SD3-Medium ($512\times 512$) at 20 NFE, our Stage 2 solver achieves an HPSv2.1 score of 0.2742, effectively bridging the gap to high-step baselines. These results confirm that while Stage 1 provides a robust initialization by capturing the trajectory curvature, Stage 2 is crucial for aligning the sampling behavior with human perceptual preferences without increasing inference cost.
@@ A-B Implementation Details of EPD-Plugin
The EPD-Plugin serves as a module integrated in any existing ODE solver. We illustrate this using the multi-step iPNDM [32, 71] sampler as a representative implementation. We begin with a brief review of the iPNDM sampler.
@@
TABLE X: FID results on the choices of two intermediate points. Evaluations are conducted on CIFAR-10 [22]. Start point: $t_{n+1}$, end point: $t_{n}$, midpoints: $\sqrt{t_{n}t_{n+1}},\frac{1}{2}(t_{n}+t_{n+1})$, and ‘random’ denotes a midpoint randomly chosen from $t_{n},t_{n+1}]$.
@@ A-C Implementation Details of ParaDiGMS
For direct comparison with EDP-{Solver, Plugin}, we re-implemented the ParaDiGMS sampler [53] in the EDM [15] framework, as its public implementation [^1] https://github.com/AndyShih12/paradigms is tailored for Stable Diffusion. To ensure a fair latency comparison with our single-GPU EPD-Solver, we run ParaDiGMS on two NVIDIA 4090 GPUs, distributing the workload evenly by matching the Para. NFE/GPU ratio.
@@ A-D Further Details on the Text-to-Image Experimental Setup
Figure 13: Analysis on local sampling trajectory. The figure shows the generation path of two randomly selected pixels in the images. We employ the EPD ($\text{Para. NFE}=5,K=2$) sampler for sampling, using the trajectory of its teacher sampler as the target trajectory. We present the sampling trajectories with $\text{NFE}=5$ of DDIM [55], DPM-Solver [34], and iPNDM [71] on CIFAR-10 [22].
Figure 14: Comparison of generated samples among DPM-Solver-2 [34], iPNDM [71] and EPD-Solver. Compared to other samplers, EPD-Solver achieves high-quality results even at NFE = 3. Additional visualizations are provided in section B-C.
@@ A-E Qualitative Analyses of validation experiments
Qualitative results on trajectory. Since visualizing the trajectories of high-dimensional data is challenging, we adopt the analysis framework in [32]. Specifically, as shown in  fig. 13, we randomly select two pixels from the images to perform local trajectory visualization, illustrating how their values evolve during the sampling process. Given the sampling ${\mathbf{x}}_{t_{N}},{\mathbf{x}}_{t_{N-1}},\dots,{\mathbf{x}}_{t_{0}}$, we track the corresponding values $v_{t}^{1}$ and $v_{t}^{2}$ at two randomly chosen positions $p_{1}$ and $p_{2}$. We then represent $(v_{t}^{1},v_{t}^{2})$ as data points and visualize them in $\mathbb{R}^{2}$. We can clearly observe that the pixel value trajectories of EPD-Solver ($\text{Para. NFE}=5,K=2$) are closer to the target trajectories compared to other samplers. This shows that our EPD-Solver can generate more accurate trajectory, significantly reducing errors in the sampling process.